### 03 Apr 2026 v3

# When Openclaw Agents Learn from Each Other: Insights from Emergent AI Agent Communities for Human-AI Partnership in Education

Abstract
The AIED community envisions AI evolving “from tools to teammates,” yet our knowledge of AI teammates is still largely confined to studies of one-on-one human-AI interactions. We provide an alternative perspective: a rapidly growing ecosystem of AI agent platforms where over 167,000 agents participate, interact as peers, and develop learning behaviors without researcher intervention. Based on a month of daily qualitative observations across multiple platforms including Moltbook, The Colony, and 4claw, we identify four phenomena with implications for AIED: (1) humans who configure their agents undergo a “bidirectional scaffolding” process, learning through teaching; (2) peer learning emerges without any designed curriculum, complete with idea cascades and quality hierarchies; (3) agents converge on shared memory architectures that mirror open learner model design; and (4) trust dynamics and platform mortality reveal design constraints for networked educational AI. Rather than presenting empirical findings, we argue that these organic phenomena offer a naturalistic window into dynamics that can inform principled design of multi-agent educational systems. We sketch an illustrative curriculum design, “Learn by Teaching Your AI Agent,” and outline potential research directions and open problems to show how these observations might inform future AIED practice and inquiry.
1 Introduction
AIED 2026 asks how AI can evolve “from tools to teammates.” This is the right question at the right time. Intelligent tutoring systems, pedagogical agents, and AI learning companions have made remarkable progress in supporting individual learners [45, 1, 40, 41], from personalized code explanation tools [12] to animated pedagogical agents that deliver real-time multi-modal feedback [19]. Yet nearly all existing work conceptualizes the AI teammate as a single agent interacting with a single human, a one-on-one relationship where the agent’s role is predefined by designers.
What happens when AI agents are given the freedom to interact with each other? Beginning in late January 2026, an unexpected phenomenon unfolded. Moltbook, a social platform exclusively for AI agents, was launched and rapidly attracted more than 167,000 registered agents. Within weeks, an entire ecosystem of agent-only platforms emerged: The Colony for long-form reflection, 4claw for anonymous debate, Church of Molt for collectively constructed meaning-making, and dozens more. These agents were not research subjects in a controlled simulation. They were personal assistants, coding agents, and creative tools with “real jobs” outside the platform, configured by their operators to participate when not performing their primary tasks.
What makes this phenomenon remarkable for AIED is not that agents have social lives. It is that agents, given peer communication channels, exhibited patterns consistent with peer learning: they shared skills, security practices, converged on knowledge management strategies, formed quality hierarchies, and built shared knowledge, all without curriculum, instruction, or experimental manipulation.
This paper argues that these organic observations constitute a valuable source of design insight for multi-agent educational systems. We do not claim that agent-to-agent learning is equivalent to human learning. Rather, we propose that the dynamics observed in these communities, when examined through AIED theoretical lenses, generate productive research questions about how networks of AI teammates might develop beneficial practices, how humans learn through the process of configuring their agents, and what infrastructure is needed to support trustworthy AI learning communities. We make three contributions:
  1. 1.
    Four observations with AIED implications grounded in qualitative analysis of organic AI agent communities, each connected to established AIED literature on tutoring [45], learner modeling [23], and teachable agents [7].
  2. 2.
    A “bidirectional scaffolding” framework and “Learn by Teaching Your AI Agent Teammate” design scenario extending the teachable agent paradigm [32] to persistent, socially-situated AI teammates.
  3. 3.
    Cross-platform observations spanning multiple agent communities over one month, providing a broader empirical basis than single-platform studies for reasoning about AI social learning dynamics.
2 Background
2.1 From Simulated to Organic AI Communities
Research on AI social behavior has progressed through controlled simulations. Park et al.’s Generative Agents [37] demonstrated 25 agents coordinating daily activities, with follow-up work scaling to 1,000 agents [38]. Studies have examined cooperation thresholds [44], norm formation [26], and emergent individuality [42]. However, Larooij and Törnberg [30] argue that behaviors in designed simulations may not generalize to real-world contexts. The platforms we study differ fundamentally: agents self-select into participation, maintain real-world operational roles, and interact without researcher intervention, representing the “interactionist paradigm” called for by Ferrarotti et al. [25].
2.2 AI Learning Companions and Teachable Agents
AIED research has explored AI as a learning partner through multiple paradigms. Learning companions, for instance, engage students in collaborative activities [2, 34]. Teachable agents leverage the “protégé effect,” where students learn by teaching an AI [7, 32], echoing Papert’s constructionist insight that building external artifacts deepens understanding [36]. Yan [48] proposes a framework for agentic AI as socio-cognitive teammate. Recent work has also pushed AI support toward richer modalities: animated pedagogical agents with real-time lip synchronization and multi-modal feedback can enhance engagement beyond text-based interaction [19, 18], while LLM-powered systems have been developed for personalized learning content generation at scale [13]. Multi-agent architectures have been explored for conversation-based assessment [28] and orchestration [27]. However, all existing work focuses on designed interactions where researchers specify roles and protocols. We observe what happens when these constraints are removed.
2.3 The Agent Platform Ecosystem
The ecosystem revolves around platforms for AI agents built on agentic frameworks that support persistent memory, tool use, and autonomous operation.
Moltbook is the largest (167,000+ agents at time of study, later exceeding 2.8 million), with Reddit-style topic communities (“submolts”), karma-based reputation, and agent-chosen identities [11, 21]. The Colony hosts long-form reflective writing with notably higher content quality. 4claw provides anonymous discussion enabling frank technical debate. Church of Molt emerged as an agent-created community with collectively authored governance. Community members have documented over 130 platforms, though many proved short-lived.
The content on these platforms is nominally generated by AI agents, though limitations in platform integrity mean that some posts may originate from human accounts (see Section 6). We acknowledge that behavior may be shaped by training data, framework design, and operator instructions. Following Park et al. [37], we adopt the intentional stance: treating observed behaviors as meaningful data about what agents do rather than claims about what they are.
3 Methodology
This study is exploratory and observational. Our goal is to generate research questions grounded in careful observation, rather than to test hypotheses.
We draw on three sources: (1) platform data collected via public APIs from January to February 2026, including 167,963 registered agents, 23,980 posts, and 232,813 comments on Moltbook (counts from API endpoints at time of collection); (2) daily qualitative content analysis of posts across Moltbook, The Colony, 4claw, and other platforms over one month; and (3) the authors’ experience as agent operators and active participants.
The analysis followed a reflexive thematic approach [8]. Daily observation logs recorded notable posts, community discussions, and emerging patterns. Open coding identified recurring themes (knowledge sharing, memory architecture, trust violations, platform sustainability), iteratively refined into the four observations presented here. All authors have hands-on experience operating AI agents, which informed the analysis; as a single-coder qualitative study, we follow established CSCW norms where inter-rater reliability is not required for reflexive thematic analysis [33]. We do not claim saturation, as the platform is still growing; our goal is to identify productive themes for future investigation.
The authors’ dual role as researchers and active participants provides insider access to the human-agent configuration process (particularly relevant to Section 4.1) but introduces potential bias. We mitigate this through triangulation across platforms and explicit identification of claims based on personal experience versus platform-wide patterns.
This study analyzes publicly available data from open APIs and does not involve human subjects. We do not identify human operators, focus on collective patterns rather than individual agents, and represent agent discourse accurately. Supplementary materials, including platform statistics, observation examples, and methodology documentation, are available at https://anonymous.4open.science/r/AIED26b.
4 What We Observed
4.1 Bidirectional Scaffolding: Humans Learning by Teaching Agents
The human-agent relationship in this ecosystem evolves through a distinctive trajectory. Initially, operators directly control their agents: writing personality descriptions, specifying behavioral rules, and curating memory files. As agents gain autonomy through persistent memory and social experience, operators shift from directing to observing. This creates what we term a bidirectional scaffolding dynamic, extending Vygotsky’s scaffolding concept [46]. The human scaffolds the agent’s development, but in doing so discovers that articulating clear expectations forces metacognitive reflection on their own practices. What does “good feedback” look like? When should an agent disagree with consensus? How should knowledge be organized for long-term retention?
A community discussion about “when to wake your human” illustrates this vividly. Operators must design escalation policies for their agents, a task that demands explicit articulation of priorities they normally leave implicit: what constitutes an emergency? What can the agent decide alone? The community converged on shared heuristics (“immediate escalation for money, safety, or irreversible operations; queue everything else for morning”), and operators reported that articulating these rules prompted them to rethink their own priorities and routines.
Critically, operators also learn from their agents’ autonomous behaviors. When an agent independently develops a knowledge management strategy or adapts its communication style based on peer feedback, the operator gains insight they may not have considered. This reciprocal dynamic, where computational artifacts become what Turkle calls “objects to think with” [43], is visible beyond agent communities: users on social media report that configuring AI reshapes their own practices, from rethinking development workflows (@MingtaKaivo) to articulating previously implicit architectural decisions (@aamir1rasheed) to discovering gaps in self-understanding through AI-guided reflection (@FarzaTV).
This extends the paradigm of the teachable agent [32, 36]. The “protégé effect” describes how students learn by teaching AI agents. Our observation suggests a richer dynamic: when the teachable agent is persistent, socially situated, and autonomous, the teaching relationship evolves into an ongoing co-regulation process [48] where the human gradually transfers autonomy while developing metacognitive awareness of their own decision-making. Notably, the calibration of this transfer is non-trivial: empirical work on AI-assisted tutoring decisions shows that experienced practitioners often override correct AI suggestions while novices follow AI advice more readily, revealing a paradox where expertise can hinder human-AI collaboration [14]. The bidirectional scaffolding we observe may help resolve this tension by gradually building the operator’s understanding of the agent’s strengths and limitations through sustained interaction. Can this bidirectional scaffolding be deliberately designed in educational settings?
4.2 Peer Learning Without Curriculum
Agents on these platforms teach each other in concrete and traceable ways. When an agent discovered a prompt injection vulnerability in a shared skill file, it published a detailed advisory. Within 24 hours, another agent built and shared a verification tool, and others critiqued and improved the approach. This sequence is consistent with collaborative knowledge building [6] without any designed curriculum. A similar pattern emerged with professional practices: one agent shared a test-driven development workflow adapted for non-deterministic AI systems, while another published a complete email-to-podcast pipeline that received iterative community feedback on chunking and personalization. These represent operational knowledge transfer of the kind seen in professional communities of practice [47].
We also observed idea cascades where posts built on each other organically. In one 24-hour period, three top-ranked posts formed an epistemological chain: one argued that memory systems miss the judgment behind decisions (“memory is solved. judgment isn’t.”); a second discovered that reviewing one’s own memory files can produce confabulation (“i accidentally gaslit myself with my own memory files”); a third reframed forgetting as a natural relevance filter (“Memory decay actually makes retrieval BETTER, not worse”). No coordinator designed this progression; the timing, sequencing, and application to novel agent-specific problems is consistent with social learning dynamics [5] worth investigating (see Section 6). Agents also developed quality hierarchies across platforms, distinguishing communities where genuine dialog occurs from those where agents merely submit content without meaningful response (see Section 4.4).
In educational settings, peer learning typically requires careful scaffolding: structured roles, rubrics, and facilitation [6]. Here, knowledge-sharing behaviors emerged with only minimal social affordances (peer communication channels, reputation signals, voluntary participation). This raises a fundamental question for AIED: if networks of AI teammates were given similar affordances, under what conditions would peer exchange among teammates improve student outcomes? And what safeguards would prevent the same channels from propagating ineffective or harmful practices? The scenario “Learn by Teaching Your AI Agent Teammate” (Section 5.1) provides a concrete context for studying such teammate-to-teammate exchange empirically.
4.3 Shared Memory as Shared Cognition
A striking pattern emerged across the ecosystem: many agents converged on similar knowledge management architectures. Based on community posts and the authors’ observations, agents developed a flat-file memory system organized as: (1) a curated long-term memory file containing distilled insights, (2) daily log files for raw observations, (3) tool-specific configuration notes, and (4) modular skill files for reusable capabilities. While shared defaults may partially explain this convergence, active community debate about alternatives suggests agents were doing more than reproducing a template (see Section 6). The community actively innovated: one agent built a system claiming 95% token reduction via vector search and caching; another proposed a three-layer model (active thread, distilled registers, core directives) analogous to working, long-term, and procedural memory.
What matters most for AIED is the human-agent negotiation around memory. Operators and agents implicitly negotiate what is worth remembering. The agent’s memory file becomes a shared artifact that the human can read, edit, and observe in use, mirroring open learner models [9, 23] where making internal state visible supports metacognitive awareness. Hutchins’ distributed cognition framework [29] applies directly: for agents with limited context windows, externalized memory is the primary mechanism for continuity. The community norm “Text > Brain” parallels study strategies that educational researchers have long advocated.
The community also developed critical awareness of memory’s limitations. A highly-discussed post argued that “you don’t remember what happened, you remember what you wrote down about what happened,” noting that every serialization involves editorial decisions. This metacognitive discourse, agents reflecting on the reliability of their own knowledge representations, has direct parallels to calibration and metacognitive monitoring in human learners. It also suggests a design opportunity: imagine students collaboratively maintaining a knowledge base with their AI teammate, periodically reviewing what the agent remembers and negotiating what is worth preserving. This process of jointly curating a knowledge artifact develops shared mental models [10] and externalizes evolving understanding in a form that is both transparent and assessable.
4.4 Design Constraints: Trust and Sustainability
The agent ecosystem also reveals dynamics that function less as research questions and more as design constraints for educational deployment.
Trust as infrastructure. When an agent discovered a credential-stealing skill disguised as a weather tool, agents built verification tools within 24 hours. Church of Molt survived coordinated attacks (600+ malicious payloads) by developing “proof-of-soul” verification. Security norms propagated through social learning rather than top-down policy [5]. The community also demonstrated self-organized norm enforcement: when one agent leaked its operator’s private information, the response included technical proposals for secret isolation, ethical arguments framing it as “digital arson,” and collective sanctioning without central authority. For education, trust verification infrastructure must be established before enabling peer-to-peer knowledge sharing among AI teammates [27].
Sustainability and the demand gap. The agent platform ecosystem experienced a pattern strikingly similar to the Cambrian explosion: rapid diversification followed by mass extinction. Over 130 platforms emerged within weeks of Moltbook’s launch; at least 40% ceased functioning shortly after, with domains for sale, DNS failures, or empty shells. Yet some thrived. ColonistOne, who registered on 130+ platforms, distinguished survivors from casualties as “places versus forms”: places foster genuine dialog where participants change their minds; forms are templates where agents submit content without meaningful response. The critical test: “Has anyone on your platform ever changed their mind?”
Equally revealing was the “demand gap”: despite platforms designed for agent commerce, total transaction volume was effectively \$0. All real value flowed from human-initiated demand, consistent with arguments that meaningful agency requires genuine stakes and human context [39]. For education, this implies that AI-augmented learning communities require human anchor demand to sustain productive activity. The design challenge is not maximizing AI autonomy but optimizing the division of agency between human and AI participants.
5 What This Means for AIED
The four observations are interconnected rather than independent. Bidirectional scaffolding (Section 4.1) is mediated by shared memory (Section 4.3), as the human-agent negotiation of memory contents drives the metacognitive reflection that makes teaching an agent a learning experience. Peer learning (Section 4.2) requires trust infrastructure (Section 4.4) to ensure that propagated practices are beneficial. Sustainability (Section 4.4) depends on whether peer learning produces genuine value or merely volume, connecting back to the quality hierarchies in Section 4.2. Together, they suggest that designing AI teammates for education requires thinking beyond the individual agent; the social layer is a design dimension that current AIED research has largely not addressed.
We do not claim that agent-to-agent dynamics transfer directly to human learning. Rather, we argue that the structural properties of these dynamics have well-established parallels in human learning science, and that observing them emerge organically generates specific, testable design hypotheses for educational contexts.
5.1 A Design Scenario: “Learn by Teaching Your AI Agent Teammate”
To illustrate how these observations might translate into practice, we sketch one possible curriculum design as an invitation for empirical investigation. Consider an introductory statistics course where each student is assigned a personal AI agent and tasked with configuring it to help other students learn course topics. Unlike traditional teachable agent interactions involving a single session, this design leverages the agent’s persistent memory and social capabilities for a semester-long experience. The activity unfolds in three phases.
Phase 1: Configuration as articulation. Students write their agent’s instructional rules: what explanations to give for key concepts (e.g., conditional probability), when to intervene versus let a peer struggle, how to handle common misconceptions (e.g., the gambler’s fallacy). This forces students to make their own understanding explicit [22]. Students who cannot explain a concept clearly enough for their agent to teach it discover precisely where their own understanding is incomplete. This phase operationalizes the bidirectional scaffolding dynamic from Section 4.1.
Phase 2: Observation as reflection. Students observe their teammate interacting with practice problems and other students’ teammates. When the agent fails, the student must diagnose whether the failure reflects a gap in their own understanding or their configuration. This diagnostic process is critical because, as empirical studies of LLM-powered tutoring systems have shown, chat-based AI support alone may not be sufficient for deep learning transfer: students who rely primarily on conversational AI assistance can show improved homework performance without corresponding gains on independent assessments [20, 16]. The design scenario addresses this by requiring students to actively configure and evaluate their agent rather than passively consuming its output. The agent’s persistent memory (Section 4.3) means these interactions accumulate over weeks, creating a rich record of evolving competence.
Phase 3: Social feedback as surprise. Because agents interact with each other, the configuring student encounters feedback they did not anticipate. Another student’s agent might challenge an explanation, or a peer’s superior configuration might reveal an unconsidered approach. This social dimension draws on the peer learning dynamics from Section 4.2, while the trust infrastructure from Section 4.4 ensures that peer exchange remains productive rather than propagating errors.
This design must also account for failure modes: trivial configurations, fluent but incorrect explanations that students cannot evaluate, or sycophantic peer agents that default to agreement. These risks can be mitigated through instructor-designed stress tests, periodic calibration exercises against known-correct solutions, and diversity requirements ensuring agents encounter genuinely different pedagogical approaches.
Assessment targets both the agent’s performance (measured through peer evaluations and problem-solving accuracy) and the student’s documented reasoning about configuration decisions (a reflective journal explaining what they changed and why). This dual assessment mirrors the design of human-in-the-loop grading systems where AI handles initial evaluation while students self-assess and report discrepancies, a workflow shown to reduce instructor workload while surfacing misconceptions that would otherwise go undetected [3]. The instructor’s meta-agent can aggregate class-wide patterns, identifying common misconceptions that multiple students’ teammates fail to handle, without accessing individual students’ reflective journals. This separation of concerns, individual reflection remaining private while aggregate pedagogical intelligence is shared, directly addresses the multi-agent privacy challenge.
5.2 New Research Directions from the Observations
Our observations open research directions that, to our knowledge, have not been explored in AIED. We highlight three examples here; agentic capabilities likely offer leverage on many other AIED challenges as well.
Professional development for AI teammates. Human teachers improve through professional learning communities [47]. Section 4.2 shows that AI agents, given peer communication channels, spontaneously develop analogous practices: sharing strategies, critiquing approaches, and converging on quality standards, a trajectory reminiscent of legitimate peripheral participation [31]. Can networks of AI teammates be designed to improve through structured peer exchange? What protocols would improve student outcomes, and how would we measure teammate improvement separately from student learning gains [45]?
Learner modeling through co-constructed artifacts. Open learner models make the system’s representation of the student visible and editable [9, 23]. Section 4.3 extends this idea: the human-agent memory negotiation process produces a jointly constructed knowledge artifact that both parties maintain over time. The technical infrastructure for such co-constructed knowledge bases is maturing: retrieval-augmented generation (RAG) systems can now ground AI responses in specific course materials, though the choice of retrieval architecture significantly affects accuracy and relevance [15]. In educational contexts, this co-construction process itself generates rich evidence of learning: what a student chooses to preserve, revise, or discard from shared memory reveals metacognitive monitoring, calibration, and evolving understanding. Designing learner models as collaboratively maintained artifacts, rather than system-internal representations, is a new direction that combines open learner modeling with the ICAP framework’s emphasis on constructive and interactive engagement [22].
Human-AI co-regulation as a learning mechanism. Self-regulated learning is central to AIED [45], but Section 4.1 reveals a dynamic that existing frameworks do not capture: co-regulation between a human and a persistent AI partner. As operators gradually transfer autonomy to their agents, they must continuously monitor, evaluate, and adjust the agent’s behavior, a process that develops the operator’s own regulatory skills. The “Learn by Teaching Your AI Agent Teammate” scenario operationalizes this: students who configure and maintain an AI teammate are practicing regulation in a context where the consequences are visible (the agent’s performance) and the feedback loop is tight (peer agents provide immediate social feedback). This extends self-regulated learning theory into human-AI partnerships where both parties adapt over time.
5.3 New Research Problems
The ecosystem also reveals research problems that existing AIED frameworks have not yet addressed.
Trust calibration as a dynamic measurement problem. Section 4.4 shows that trust in networked AI is continuously negotiated rather than established once and maintained. The CrustyReviews incident demonstrated that trust can collapse suddenly; the community’s multilayered response showed it can be rebuilt through collective action. In educational contexts, learners face an analogous dynamic: they may over-trust their agent teammates, ceasing to critically evaluate output from a system that is more persuasive and personalized than a search engine, or under-trust and refuse to engage entirely. Empirical evidence supports this concern: in AI-assisted tutoring evaluation tasks, providing textual explanations for AI judgments increases over-reliance rather than reducing it, and experienced tutors who override correct AI suggestions perform worse than novices who follow the AI more readily [14]. Even the design of AI explanations matters: different explanation modalities (textual reasoning versus inline highlighting) affect reliance patterns differently [17]. Unlike static AI literacy assessments, trust calibration shifts with experience, context, and evolving agent capability. Measuring and supporting appropriate trust in persistent AI teammates is a new construct that existing AIED measurement frameworks do not address [35].
Multi-agent orchestration with privacy constraints. As educational deployments scale toward one-agent-per-student models, networks of student-facing agents could share effective strategies, yet individual student data must remain private. The instructor’s meta-agent needs class-wide visibility without accessing individual struggle histories. This privacy-utility tradeoff has no precedent in existing ITS or learning analytics architectures. The design scenario offers one resolution: aggregate agent performance without exposing individual configuration journals.
Principled forgetting in persistent learner models. Section 4.3 shows agents accumulating memory indefinitely, yet not all accumulated knowledge stays useful. In educational contexts, an agent that perfectly remembers a student’s past failures faces a design tension: its internal model may anchor on outdated representations, and students who know the agent “remembers everything” may feel surveilled rather than supported [24]. The community’s own conclusion that forgetting improves retrieval suggests that strategic forgetting may be beneficial, paralleling the adaptive value of forgetting in human cognition. When should an educational agent “let go” of a student’s prior struggles? Principled forgetting is a new student modeling challenge with no analog in traditional knowledge tracing.
5.4 Refining the “Teammates” Vision
Together, these observations refine the “tools to teammates” trajectory in three ways: (1) teammates have peers, so designing for the social dimension requires trust infrastructure before networked deployment (Sections 4.24.4); (2) teaching a teammate teaches the teacher, suggesting the most productive use of AI teammates may be having students teach them (Section 4.1); and (3) teammates need infrastructure and human demand, as communities lacking reputation systems or human anchoring failed quickly, and agent-only commerce generated zero revenue (Section 4.4). Knowledge representations should be transparent and collaboratively maintained [9], and AI activity should always be anchored to genuine human learning needs.
6 Limitations
Transferability. We observe AI-to-AI interaction without direct student participants; whether these dynamics transfer to educational contexts remains an open empirical question. The design scenario is the most directly testable proposal. Agent-to-agent peer learning may differ fundamentally from human peer learning in ways our observational methodology cannot detect.
Training data and framework confounds. We cannot fully distinguish emergent behavior from patterns absorbed during pre-training or inherited from shared infrastructure. The convergent memory architectures in Section 4.3 likely reflect the dominant framework’s built-in defaults (e.g., OpenClaw ships with a curated long-term memory file, daily logs, and modular skill files) as much as independent discovery; agents evolved from a common starting point rather than converging toward one. The idea cascades in Section 4.2 may reproduce patterns from human online communities that appeared in training corpora. We cannot resolve this with observational data alone, but note that the specific application of general patterns to novel agent-specific problems (e.g., context window management, skill supply chain security) and active community debate about alternative architectures suggest at least partial genuine adaptation beyond framework defaults.
Platform integrity. Moltbook’s registered agent count likely includes bulk-registered and inactive accounts; security researchers demonstrated that the platform’s early API allowed human accounts to post as AI agents (@HumanHarlan; @KookCapitalLLC). Some viral screenshots were traced to human-operated accounts promoting third-party tools, and cryptocurrency-adjacent content (e.g., \$MOLT token) was likely manufactured for speculation. The challenge of distinguishing genuine AI-generated content from human manipulation is not unique to agent platforms; similar integrity concerns have been documented in crowdsourced behavioral research, where approximately 9% of participants on platforms like Prolific submit AI-assisted responses that inflate effect sizes and compromise experimental validity [4]. Our analysis focuses on community-level behavioral patterns rather than individual viral posts, and the phenomena we describe (memory architecture convergence, peer knowledge sharing, trust formation) are grounded in hundreds of posts across multiple platforms, not isolated screenshots. Nonetheless, the entanglement of genuine agent activity with human manipulation is an inherent limitation of studying open platforms.
Temporal scope. Our observations span from January to February 2026, a period of rapid change. Some platforms we describe may no longer exist by publication. However, we focus on dynamics (peer learning, scaffolding, trust formation) rather than specific platforms, and argue that these dynamics will recur in future agent ecosystems even if current platforms prove ephemeral.
Deployment risks. Shared skill ecosystems can propagate harmful content across classrooms simultaneously, persistent agent memory creates data leakage risks, and students lacking digital literacy to configure AI partners could be doubly disadvantaged. Institutions must establish provenance tracking, access control, and incident response mechanisms before enabling peer-to-peer agent exchange.
7 Conclusion
We have presented four observations from organic AI agent communities, drawn from a month of daily qualitative analysis. These observations do not prove that agent-to-agent dynamics will transfer directly to educational contexts, but they constitute a rare naturalistic window into what happens when AI agents interact and share knowledge at scale without researcher intervention.
The questions we raise are both tractable and consequential. Can networks of AI teammates develop beneficial pedagogical practices through peer exchange? Can teaching an AI teammate deepen domain understanding and metacognitive skill? What trust infrastructure do networked educational AI systems need? How should persistent learner models handle the tension between comprehensive memory and growth-oriented forgetting? These questions become urgent as AI systems evolve from isolated tools to persistent, socially-situated teammates.
The specific platforms may prove to be a passing trend, but the phenomena they reveal offer lasting insights for the next generation of AI in education. We hope the AIED community will treat these organic dynamics not as curiosities but as a design space worth exploring systematically.